# Pulling apart the mechanisms that lead to jammed knitted fabrics

### v1

@@ Acknowledgements.
For the contact force, we repeat this analysis on the z-component of the same row contact for a variety of yarn lengths with a scaling compression contant of $k=0.1mN/mm^{2}$. This contact force is relatively linear and consistent for both jammed and unjammed fabrics. These simulations were stretched in the y-direction and all have the same step size, $\Delta a_{y}=0.001mm$, and each RMS was calculated for the first 20 force vs strain points, displayed in 5. The average RMS over all yarn lengths is $2.24*10^{-5}$ N.
@@
$k$ [mN/mm2] Slope [J/mm] Y-Intercept [J] Max Abs(Residual) [J] Mean Abs(Residual) [J] MSE [J]
@@
Table 6: Fitting information for the linear fits to the jammed regime of the contact energy gap data for various values of the compression scaling constant $k$. The maximum value of the absolute value of the residuals and the mean of the absolute value of the residuals are listed. The mean absolute value of the residuals is within our estimated energy gap error for all but one linear fit ($k=1.1mN/mm^{2}$). The mean squared error is within our estimated energy gap error for all but five values of $k$ ($k={0.8,1.0,1.1,1.5,2.0}$).
@@
Figure S15: Energy gap for the contact (pink) and bending (blue) energies as a function of length of yarn per stitch for compression scaling constant $k=0.1$ mN/mm2 pulled in the x-direction (a) and the y-direction (b) on a log scale. This is a re-scaling of Main Text Fig. 3. All simulations where the energy gap is zero are not present on this graph due to the logarithmic scaling. The light green region shows where in length of yarn per stitch parameter space the fabric moves from the jammed transition region ($\Delta E_{c}>5*10^{-5}$ J) to un-jammed ($\Delta E_{c}<5*10^{-5}$ J). We can see a change in behavior in the jammed energy gap when $L/r$ approaches the green end-of-transition region, supporting our error floor of $5*10^{-5}$. (c,d) Energy gap rescaled by the total energy of the rest, zero-force configuration in the (c) x- and (d) y-directions. This provides an intrinsic view of the extrinsic definition of the energy gap (Main Text Eq. 4).

### v2

@@ Acknowledgements.
For the contact force, we repeat this analysis on the $z$-component of the same-row contact for a variety of yarn lengths with a scaling compression constant of $k=0.1~{}{\rm mN/mm^{2}}$. This contact force is relatively linear and consistent for both jammed and unjammed fabrics. These simulations were stretched in the $y$-direction and all have the same step size, $\Delta a_{y}=0.001~{}{\rm mm}$, and each RMS was calculated for the first 20 force vs strain points, displayed in 5. The average RMS over all yarn lengths is $2.24*10^{-5}$ N.
@@
$k$ [mN/mm2] Slope [J/mm] Y-Intercept [J] Max Abs(Residual) [J] Mean Abs(Residual) [J] MSE [J2]
@@
Table 6: Fitting information for the linear fits to the jammed regime of the contact energy gap data for various values of the compression scaling constant $k$. The maximum value of the absolute value of the residuals and the mean of the absolute value of the residuals are listed. The mean absolute value of the residuals is within our estimated energy gap error for all but one linear fit ($k=1.1{\rm mN/mm^{2}}$). The mean squared error is within our estimated energy gap error for all but five values of $k$ ($k={0.8,1.0,1.1,1.5,2.0}$).
@@
Figure S15: Energy gap for the contact (coral) and bending (blue) energies as a function of length of yarn per stitch for compression scaling constant $k=0.1$ mN/mm2 pulled in the $x$-direction (a) and the $y$-direction (b) on a log scale. This is a re-scaling of Main Text Fig. 3. All simulations where the energy gap is zero are not present on this graph due to the logarithmic scaling. The light green region shows where in length of yarn per stitch parameter space the fabric moves from the jammed transition region ($\Delta E_{c}>5*10^{-5}$ J) to un-jammed ($\Delta E_{c}<5*10^{-5}$ J). We can see a change in behavior in the jammed energy gap when $L/r$ approaches the green end-of-transition region, supporting our error floor of $5*10^{-5}$ J. (c,d) Energy gap rescaled by the total energy of the rest, zero-force configuration in the (c) $x$- and (d) $y$-directions. This provides an intrinsic view of the extrinsic definition of the energy gap (Main Text Eq. 4).